Rony will inherit $ 250000 his grandmother in 5 year hence he decided to take a loan now for buy a house for his daughter that will cost 500000 dollars now what amount should he borrow and what will the percentage of down payment pay for the mortgage of house


1
Expert's answer
2021-05-24T19:16:48-0400

A=500000

P= 25000

T = 5 years

A= p+(PRT/100)

A-P = PRT/100

Therefore,

(500000-250000) = (250000*R * 5) /100

(250000*100) = 1250000R

Hence; R= 25000000/1250000

R= 20%